OpenStudy (anonymous):

Solve the equation 3x2+24=0 for x

thomaster (thomaster):

\(\large\sf 3x^2+24=0\) get the numbers to the right, and the x's to the left. First step would be to subtract 24 from both sides.

thomaster (thomaster):

then divide the whole thing by 3

thomaster (thomaster):

then you'll get \(\large\sf x^2=-8\) x would be \(-\sqrt{8}\)